OpenStudy (anonymous):

Hi. I appreciate any help I can get on this question. Thanks in advance. A calorimeter contains 500 g of water at 25°C. You place a hand warmer containing 100g of liquid sodium acetate (NaAC) inside the calorimeter. When the sodium acetate finishes crystallizing, the temperature of the water inside the calorimeter is 32.2°C. The specific heat of water is 4.18 J/g-°C. What is the enthalpy of fusion (ΔHf) of the sodium acetate? Show your work.

ganeshie8 (ganeshie8):

@chmvijay

OpenStudy (anonymous):

I have this part but I am not sure what to do next. q = mCΔT q = 500 * 4.18 * (32.2 - 25) q = 500 * 4.18 * 7.2 q = 15,048 J

OpenStudy (wolfe8):

If you did that right, I believe that will be the energy used to crystalize the NaAC. Enthalpy of fusion is the energy required to crystalize every mole of the substance. I might have forgotten this so heads up if it's wrong.

OpenStudy (kewlgeek555):

@soccersoccer1 : \(\Huge\bf Welcome~to~ \color{#00B4ff}{Open}\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge\bf\color{#7cc517}{Study~!!}~\) I am not that advanced into chemistry yet, so I can't help you. Sorry. But @wolfe8 seems to have it handled here. ;) Good luck, have fun, keep learning, get A's, ~Ezra (kewlgeek555)

OpenStudy (kainui):

\[q=m*\Delta H_f\] So this says the total heat is equal to the mass times the heat of enthalpy, just divide q by m to get the enthalpy of fusion! Think of the enthalpy of fusion as being the amount of heat you need per mass and it's for the most part a constant.

OpenStudy (kainui):

So for instance if I wanted to freeze water into ice it would take the same ratio of heat to mass to freeze it. The more water you want to freeze, the more heat you have to lose right? But that amount will be constant, and that's your enthalpy of fusion.

OpenStudy (anonymous):

So the enthalpy of fusion would be 15,048/500 = 30.096? What is the 100 g in the question?

OpenStudy (kainui):

Ahh, no, the mass of the water heated up is used to calculate Q, but to calculate delta H you need the mass of the NaAC itself.

OpenStudy (kainui):

So enthalpy of fusion would be 15,048J/100g if that makes sense.

OpenStudy (anonymous):

Oh so 100 g is the mass of the NaAC.

OpenStudy (kainui):

First you were calculating how much heat was transferred from the NaAC to the water. Then you found out how much of that heat came per mass of the NaAC itself. Exactly. =)

OpenStudy (anonymous):

Got it. Thanks so much for your help. :)